OpenStudy (anonymous):

How do I factor 16x^2 -49

OpenStudy (haseeb96):

did u not get my point ?

OpenStudy (anonymous):

Its not what I was looking for

OpenStudy (anonymous):

@haseeb96

OpenStudy (haseeb96):

u are asking that how can u factor it?

OpenStudy (anonymous):

It has something to do with this a^2-b^2 ➡(a-b)(a+b)

OpenStudy (haseeb96):

16x^2-49 for factorization take this equation equal to zero

OpenStudy (haseeb96):

do u know this formula a^2 - b^2 = (a-b)(a+b)

OpenStudy (anonymous):

That's the one in talking about

OpenStudy (haseeb96):

if u know then have a look on the equation all the values in the equation are perfect square

OpenStudy (haseeb96):

haha okay

OpenStudy (haseeb96):

4^2=16 7^2=49 we can write 16x^2-49 = (4x-7)(4x+7) (4x-7)(4x+7) this is your answer

OpenStudy (anonymous):

Thank you so much man

OpenStudy (haseeb96):

and most welcome lady :)

OpenStudy (anonymous):

Lol